OpenStudy (anonymous):

I WILL MEDAL AND FAN JUST HELP ALREADY!!!!! I just need help with this: Part A: Write the expression x2 + 7x + 12 as a product of two linear expressions. Show your work and justify each step. Part B: Rewrite x2 + 4x + 4 as a square of a linear expression. Part C: Do the expressions in parts A and B have a common factor? Justify your answer.

OpenStudy (anonymous):

@triciaal @mathmale @mathstudent55 @zepdrix @pooja195 @ParthKohli @jim_thompson5910 @Opcode

OpenStudy (khally92):

I think you should solve the quadratic equations for boths a and b

OpenStudy (khally92):

a) gives you \[(x+3)(x+4)\] b) gives you \[(x+2)(x+2)=(x+2)^2\]

OpenStudy (anonymous):

@khally92 How did you get your answer for b?

OpenStudy (anonymous):

@triciaal Help

OpenStudy (anonymous):

@Zela101

OpenStudy (anonymous):

@ayeshaafzal221 @magepker728 @zepdrix

OpenStudy (md152727):

@Thebossofme191 @OceanChild @Opcode @PizzaLover123 @Agl202 @Anaise

zepdrix (zepdrix):

:o

OpenStudy (anonymous):

(I have already written down the answer for part A, I need help with Part B and C)

OpenStudy (anonymous):

How did @khally92 get (x + 2)^2 as her answer?

zepdrix (zepdrix):

\[\large\rm x^2+4x+4\]You can do factor by grouping if you've gotten good with that method.

zepdrix (zepdrix):

Factors of 4: 2*2 is one way to do it, yes?

OpenStudy (anonymous):

Do you mean x^2 + 3x + 4x + 4?

OpenStudy (anonymous):

Wait, nvm. I thot we were using the same equation as part A nvm

zepdrix (zepdrix):

:P

OpenStudy (anonymous):

Yes 2 and 2 would be the best

OpenStudy (anonymous):

x2 + 2x + 2x + 4

OpenStudy (anonymous):

(x2 + 2x) + (2x + 3)

OpenStudy (anonymous):

*(2x + 4)

OpenStudy (anonymous):

GCF Binomial 1: x Binomial 2: 2 Divide GCF x(x + 2) + 2(x + 2)

OpenStudy (anonymous):

Join together: (x + 2)(x + 2)

OpenStudy (anonymous):

OHHHHHHH

zepdrix (zepdrix):

grouping is a little bit longer of a way to do it, but ya it always works so that's nice :)

zepdrix (zepdrix):

A: (x+3)(x+4) B: (x+2)(x+2) Common factor?

OpenStudy (anonymous):

Does it start with an x and end with an x?

zepdrix (zepdrix):

Hint, this is a factor in orange here \(\large\rm \color{orangered}{(x+3)}(x+4)\)

OpenStudy (anonymous):

Ima assume that the comon factor is x, correct?

zepdrix (zepdrix):

Example: \(\large\rm x(x-2)(x+3)\) In this example, x is a factor. In your problems, both A and B, x is `not` a factor. x is `part of` a factor, but it's not a factor by itself.

OpenStudy (anonymous):

How is x + 3 a factor though?

zepdrix (zepdrix):

21 = 7*3 Therefore 7 is a factor of 21. x^2+4x+4 = (x+2)*(x+2) Therefore (x+2) is a factor of x^2+4x+4

OpenStudy (anonymous):

But how is x + 3 a common factor?

zepdrix (zepdrix):

Things that multiply together are factors.

zepdrix (zepdrix):

(x+3) is not a common factor, it's a factor of the polynomial in A, but not a factor of the polynomial in B.

OpenStudy (anonymous):

So there are no common factors then, correct?

zepdrix (zepdrix):

correct :)

OpenStudy (anonymous):

Hooray! Thanks for all your help @zepdrix . I'd give you owlbucks but I don't have any. :(

zepdrix (zepdrix):

owl what? 0_o oh those still exist? hah

OpenStudy (anonymous):

You can get letters of recommendation and stuff with them

OpenStudy (anonymous):

Anyways, ima go finish the question. ill be back in like 7 minutes. Or I won't be back at all. Either Way! :)